# Spectral Analysis and Decay Mechanisms of $1^{-+}$ Hybrid States in Light Meson Sector

### v2

@@ Abstract
The exploration of exotic mesons, which transcend the conventional quark-antiquark framework, is pivotal for advancing our understanding of QCD and the strong interaction. Among these, states possessing the quantum numbers $J^{PC}=1^{-+}$, such as $\pi_{1}(1600)$, $\pi_{1}(2015)$, and the recently discovered $\eta_{1}(1855)$, have attracted significant attention due to their potential hybrid nature, which involves gluonic excitations. However, the physical interpretation of these states remains contentious, primarily due to inconsistencies between theoretical predictions and experimental observations regarding their decay widths and mass spectra. To address these challenges, we employ a potential model inspired by SU$(3)$ lattice gauge theory to calculate the masses of light-flavor $1^{-+}$ hybrid states and utilize the constituent gluon model to analyze their strong decay properties. Our mass calculations suggest that while the $\pi_{1}(1600)$ and $\eta_{1}(1855)$ masses are in close agreement with the predicted hybrid states, the corresponding decay widths for $\pi_{1}(1600)$ and $\pi_{1}(2015)$ do not support their classification as hybrids, implying alternative interpretations, such as tetraquark or molecular configurations. Conversely, the $\eta_{1}(1855)$ is consistent with a mixed hybrid state composed of $(u\bar{u}+d\bar{d})g/\sqrt{2}$ and $s\bar{s}g$ components, with mixing angles ranging from $21^{\circ}$ to $74^{\circ}$. Additionally, we predict the masses and decay widths of yet-unobserved isoscalar and excited hybrid states, providing valuable targets for future experimental searches. This study not only clarifies the hybrid nature of certain exotic mesons but also contributes to the systematic classification of hybrid states within the meson nonet, thereby enhancing our comprehension of exotic hadronic states.
@@ I Introduction
The exploration of exotic mesons, which go beyond the conventional quark-antiquark framework, is crucial for advancing our understanding of QCD and the strong interaction. These mesons, characterized by unconventional quantum numbers, challenge our current theories and provide a unique opportunity to study the role of gluonic excitations. Among them, states with $J^{PC}=1^{-+}$, such as $\pi_{1}(1600)$ E852:1998mbq , $\pi_{1}(2015)$ E852:2004gpn and the recently observed $\eta_{1}(1855)$ BESIII:2022riz ; BESIII:2022iwi , have drawn significant attention due to their potential hybrid nature. However, interpreting the properties of these states remains a contentious issue, largely due to discrepancies between theoretical predictions and experimental observations in terms of their decay widths and mass spectra.
Several collaborations IHEP-Brussels-LosAlamos-AnnecyLAPP:1988iqi ; Aoyagi:1993kn ; E852:1997gvf ; VES:2001rwn ; CrystalBarrel:1998cfz have reported the observation of $\pi_{1}(1400)$, but its existence as a resonance remains controversial, as discussed in Meyer:2010ku . The $\pi_{1}(1600)$ has also been observed in various decay channels, including $\eta^{\prime}\pi$ Baker:2003jh , $\rho\pi$ Khokhlov:2000tk , $f_{1}(1285)\pi$ COMPASS:2009xrl , and $b_{1}(1235)\pi$ CLEO:2011upl . Over the years, many theorists have interpreted $\pi_{1}(1600)$ as a hybrid Narison:1999hg ; Shastry:2022mhk ; Chen:2023ukh ; Iddir:1988jd ; Benhamida:2019nfx ; Eshraim:2020ucw , though this paper presents an alternative viewpoint. Additionally, $\pi_{1}(1600)$ has been considered a candidate for a tetraquark state Chen:2008qw , given that gluons in a hybrid state can easily split into a quark-antiquark pair ($q\bar{q}$). Furthermore, some theories suggest that $\pi_{1}(1600)$ may be a mixing state of a molecular state and a tetraquark or hybrid Narison:2009vj . The $\pi_{1}(2015)$, observed exclusively in $b_{1}(1235)\pi$ E852:2004rfa and $f_{1}(1285)\pi$ E852:2004gpn by the BNL E852 experiment, remains poorly understood due to a lack of comprehensive experimental data. Lattice QCD calculations of radial excited states identify $\pi_{1}(2015)$ as a promising candidate for the first excited state of a hybrid Dudek:2010wm , though this contradicts the conclusions of the present work.
@@
The limited number of discovered exotic particles and the scarcity of available data present significant challenges in accurately classifying these states. To address these challenges, we adopt a potential model inspired by SU(3) lattice gauge theory, which has demonstrated success in capturing gluonic contributions to hadron interactions, to calculate the masses of light-flavor $1^{-+}$ hybrid mesons. Additionally, the constituent gluon model is employed to analyze their strong decay properties, offering a robust framework to investigate the interplay of gluonic and quark-antiquark dynamics in hybrid mesons. We also predict the masses and decay widths of yet-undiscovered isoscalar and excited hybrid states, which provide valuable targets for experimental exploration. This study not only clarifies the hybrid nature of specific exotic mesons but also enhances the systematic classification of hybrid states within the meson nonet. Furthermore, conventional mesons are known to follow Regge trajectories Anisovich:2000kxa . Investigating whether hybrid states also adhere to this pattern is crucial for establishing a complete SU(3) multiplet, as discussed by other authors Chen:2023ukh ; Yan:2023vbh ; Qiu:2022ktc . By addressing these unresolved issues, our work contributes to a deeper understanding of the internal structure of exotic hadronic states and the dynamics of the strong interaction.
@@ III Decay model of the hybrid state
The decay properties of hybrid states are essential for understanding their structure and classification. Several theoretical models have been developed to investigate the decay widths of hybrid states, including the flux tube model Page:1998gz ; Isgur:1985vy ; Close:1994hc ; Barnes:1995hc , the constituent gluon model LeYaouanc:1984gh ; Iddir:1988jd ; Ishida:1991mx ; Kalashnikova:1993xb ; Swanson:1997wy ; Iddir:2000yb ; Ding:2006ya ; Iddir:2007dq ; Benhamida:2019nfx ; Farina:2020slb ; Tanimoto:1982eh ; Tanimoto:1982wy , QCD sum rules DeViron:1984svx ; Zhu:1998sv ; Zhang:2002id ; Chen:2010ic ; Huang:2010dc ; Huang:2016upt , and lattice QCD McNeile:2002az ; McNeile:2006bz ; Woss:2020ayi . In this work, we adopt the constituent gluon model, where the hybrid meson is represented as a quark-antiquark pair coupled to a transverse electric (TE) gluon ($J_{g}^{PC}=1^{+-}$) Chen:2023ukh . This gluon acts as a dynamic degree of freedom and can annihilate into a new quark-antiquark pair, which then combines with the original pair to form two mesons. This mechanism describes how hybrid states decay into two mesons, as illustrated in Fig. 2.
@@
Selection rules play a significant role in hybrid decays. For instance, the spin selection rule prohibits decay into final states composed entirely of spin-singlet mesons if the hybrid’s $q\bar{q}$ configuration is in a spin-singlet state. Additionally, spatial overlap integrals, which describe the overlap of wavefunctions between initial and final states, generally suppress the decay widths of excited states compared to ground states.
Another decay selection rule indicates that, assuming $\beta_{B}=\beta_{C}$, the decay of the GE state into $S+S$ is forbidden Iddir:2000yb . This rule arises from the treatment of integrals under the condition $\beta_{B}=\beta_{C}$. In Ref. Iddir:1988jd , the integrals are evaluated with $\beta_{B}\neq\beta_{C}$, and the results indicate that the spatial overlap is generally proportional to $(\beta_{B}-\beta_{C})^{2}$, which is typically small. The same outcome is observed in the flux tube model Close:1994hc . Similarly, if the two S-wave states possess different internal structures or sizes, as discussed in Ref. Close:1994pr , the selection rule prohibiting $S+S$ as final states no longer applies.
@@ IV.1 The $I^{G}J^{PC}=1^{-}1^{-+}$ state
The combined analysis of masses and decay widths provides essential insights into the nature of exotic mesons, particularly those with $J^{PC}=1^{-+}$. Section II presents mass calculations for different flavor configurations, showing that the $q\bar{q}g$, $(u\bar{u}-d\bar{d})g/\sqrt{2}$, and $(u\bar{u}+d\bar{d})g/\sqrt{2}$ states are mass-degenerate. Given the quantum numbers $I^{G}J^{PC}=1^{-}1^{-+}$, we assume that $\pi_{1}(1600)$ corresponds to the configuration $(u\bar{u}-d\bar{d})g/\sqrt{2}$ in our calculations. For the decay channels involving $f_{1}(1285)$ and $f_{1}(1420)$, we model them as follows:
@@
Viewing $\pi_{1}(1600)$ as a hybrid state, we calculate its total decay width to be 15.15 MeV, with $b_{1}(1235)\pi$ identified as the primary decay channel (see Table 3). Reference Page:1998gz also reports a similar narrow decay width. However, this theoretical result exhibits a substantial deviation from the experimental values and averages reported by the PDG ParticleDataGroup:2024cfk . In our analysis, the decay width to $b_{1}(1235)\pi$ via $D$-wave is 0 MeV, implying that $\pi_{1}(1600)$ can only decay into $b_{1}(1235)\pi$ through $S$-wave. This finding contradicts the PDG’s reported branching ratio $\mathcal{B}\left(\left(b_{1}\pi\right)_{\text{D-wave}}\right)/\mathcal{B}\left(\left(b_{1}\pi\right)_{\text{S-wave}}\right)=0.3\pm 0.1$ ParticleDataGroup:2024cfk . This discrepancy raises questions about its classification as a hybrid state and points to alternative interpretations, such as tetraquark or molecular configurations.
@@
14.1 1.05 15.15 $370^{+50}_{-60}$
Similarly, $\pi_{1}(2015)$, although its mass matches lattice QCD predictions for the first excited hybrid state, exhibits narrow decay widths inconsistent with experimental observations, suggesting it may belong to a different category of exotic mesons. With the same exotic quantum numbers $I^{G}J^{PC}=1^{-}1^{-+}$, the mass of $\pi_{1}(2015)$ is reported to be $2014\pm 20^{+16}_{-16}$ MeV ParticleDataGroup:2024cfk , which is in close proximity to 2193 MeV (see Table 1). As suggested by Meyer:2010ku and previous lattice QCD (LQCD) mass calculations Dudek:2010wm , this approximate mass motivates us to propose that the $\pi_{1}(2015)$ state could be the first excited state of the hybrid $1^{-}1^{-+}$. Consequently, we assume that the $\pi_{1}(2015)$ state has the flavor configuration $(u\bar{u}-d\bar{d})g/\sqrt{2}$ and proceed with the calculation of its decay width. For the decay channels involving $K_{1}(1270)K$ and $K_{1}(1400)K$, we treat them as linear combinations of ${}^{1}P_{1}$ and ${}^{3}P_{1}$ states Page:1998gz ; Kokoski:1985is ,
@@
$\alpha_{i}=78.7^{\circ}$, see Ref. Wang:2019qyy .
As shown in Table 4, the total decay width of $\pi_{1}(2015)$ that we calculated is 0.475 MeV, which is inconsistent with the experimental values ParticleDataGroup:2024cfk . The primary decay channel of $\pi_{1}(2015)$ is $b_{1}(1235)\pi$, while other decay channels can be neglected. Experimentally, $\pi_{1}(2015)$ has been observed in the $b_{1}(1235)\pi$ channel. However, the narrow decay width leads us to question the interpretation of $\pi_{1}(2015)$ as the first excited state of the $1^{-}1^{-+}$ hybrid. Until now, $\pi_{1}(2015)$ has been observed only in the $b_{1}(1235)\pi$ and $f_{1}(1285)\pi$ channels E852:2004rfa ; E852:2004gpn .
@@
$f_{1}(1285)\pi$ $h_{1}(1170)\rho$ $b_{1}(1235)\omega$ $b_{1}(1235)\pi$ $K_{1}(1270)K$
0.03 $\approx 0$ $\approx 0$ 0.42 0.015
$K_{1}(1400)K$ $a_{1}(1260)\rho$
$\approx 0$ 0.01
@@
Furthermore, we predict the properties of the first excited isoscalar hybrids $\pi_{1}(2193)$ with $I^{G}J^{PC}=1^{-}1^{-+}$. Its decay properties are listed in Table 5. This state is expected to have narrow decay widths and distinctive dominant decay channels, such as $K_{1}(1650)K$, and $b_{1}(1235)\pi$. For the decay channel involving $K_{1}(1650)K$, we consider $K_{1}(1650)$ to be a linear combination of the $2^{1}P_{1}$ and $2^{3}P_{1}$ states, based on the mass predictions of mesons in Ref. Godfrey:1985xj , and assume it adopts the ideal mixing angle. Following the approach in Ref. Blundell:1996as , we assume a mixing angle of $45^{\circ}$ for $K_{1}(1650)^{+}$ and $K_{1}(1650)^{0}$, and $-45^{\circ}$ for $K_{1}(1650)^{-}$ and $\bar{K}_{1}(1650)^{0}$. The narrow width make these states challenging to observe experimentally, but their detection would provide valuable confirmation of the theoretical framework for hybrid mesons.
@@
$f_{1}(1285)\pi$ $K_{1}(1650)K$ $b_{1}(1235)\pi$
$0.13$ $5.15$ 1.36
$\rho(1450)\pi$ $\eta_{2}(1645)\pi$
$0.2$ $0.2$

### v3

@@ Abstract
The exploration of exotic mesons, which transcend the conventional quark-antiquark framework, is pivotal for advancing our understanding of QCD and the strong interaction. Among these, states possessing the quantum numbers $J^{PC}=1^{-+}$, such as $\pi_{1}(1600)$, $\pi_{1}(2015)$, and the recently discovered $\eta_{1}(1855)$, have attracted significant attention due to their potential hybrid nature, which involves gluonic excitations. However, the physical interpretation of these states remains contentious, primarily due to inconsistencies between theoretical predictions and experimental observations regarding their decay widths and mass spectra. To address these challenges, we employ a potential model inspired by SU$(3)$ lattice gauge theory to calculate the masses of light-flavor $1^{-+}$ hybrid states and utilize the constituent gluon model to analyze their strong decay properties. Our mass calculations suggest that while the $\pi_{1}(1600)$ and $\eta_{1}(1855)$ masses are in close agreement with the predicted hybrid states, the corresponding decay widths for $\pi_{1}(1600)$ and $\pi_{1}(2015)$ do not support their classification as hybrids, implying alternative interpretations, such as tetraquark or molecular configurations. Conversely, the $\eta_{1}(1855)$ is consistent with a mixed hybrid state composed of $(u\bar{u}+d\bar{d})g/\sqrt{2}$ and $s\bar{s}g$ components, with mixing angles ranging from $17.7^{\circ}$ to $84.2^{\circ}$. Additionally, we predict the masses and decay widths of yet-unobserved isoscalar and excited hybrid states, providing valuable targets for future experimental searches. This study not only clarifies the hybrid nature of certain exotic mesons but also contributes to the systematic classification of hybrid states within the meson nonet, thereby enhancing our comprehension of exotic hadronic states.
@@ I Introduction
The exploration of exotic mesons, which go beyond the conventional quark-antiquark framework, is crucial for advancing our understanding of QCD and the strong interaction. These mesons, characterized by unconventional quantum numbers, challenge our current theories and provide a unique opportunity to study the role of gluonic excitations. Among them, states with $J^{PC}=1^{-+}$, such as $\pi_{1}(1600)$ E852:1998mbq , $\pi_{1}(2015)$ E852:2004gpn and the recently observed $\eta_{1}(1855)$ BESIII:2022riz ; BESIII:2022iwi , have drawn significant attention due to their potential hybrid nature. However, interpreting the properties of these states remains a contentious issue, largely due to discrepancies between theoretical predictions and experimental observations.
Several collaborations IHEP-Brussels-LosAlamos-AnnecyLAPP:1988iqi ; Aoyagi:1993kn ; E852:1997gvf ; VES:2001rwn ; CrystalBarrel:1998cfz have reported the observation of $\pi_{1}(1400)$, but its existence as a resonance remains controversial Meyer:2010ku . The $\pi_{1}(1600)$ has also been observed in various decay channels, including $\eta^{\prime}\pi$ Baker:2003jh , $\rho\pi$ Khokhlov:2000tk , $f_{1}(1285)\pi$ COMPASS:2009xrl , and $b_{1}(1235)\pi$ CLEO:2011upl . Over the years, many theorists have interpreted $\pi_{1}(1600)$ as a hybrid Narison:1999hg ; Shastry:2022mhk ; Chen:2023ukh ; Iddir:1988jd ; Benhamida:2019nfx ; Eshraim:2020ucw . Additionally, $\pi_{1}(1600)$ has been considered a candidate for a tetraquark state Chen:2008qw , given that gluons in a hybrid state can easily split into a quark-antiquark pair ($q\bar{q}$). Furthermore, some theories suggest that $\pi_{1}(1600)$ may be a mixing state of a molecular state and a tetraquark or hybrid Narison:2009vj . The $\pi_{1}(2015)$, observed exclusively in $b_{1}(1235)\pi$ E852:2004rfa and $f_{1}(1285)\pi$ E852:2004gpn by the BNL E852 experiment, remains poorly understood due to a lack of comprehensive experimental data. Lattice QCD calculations of radial excited states identify $\pi_{1}(2015)$ as a promising candidate for the first excited state of a hybrid Dudek:2010wm .
@@
The limited number of discovered exotic particles and the scarcity of available data present significant challenges in accurately classifying these states. To address these challenges, we adopt a potential model inspired by SU(3) lattice gauge theory, which has demonstrated success in capturing gluonic contributions to hadron interactions, to calculate the masses of light-flavor $1^{-+}$ hybrid mesons. Additionally, the constituent gluon model is employed to analyze their strong decay properties, offering a framework to investigate the interplay of gluonic and quark-antiquark dynamics in hybrid mesons. We also predict the masses and decay widths of yet undiscovered isoscalar and excited hybrid states, which provide valuable targets for experimental exploration. This study not only clarifies the hybrid nature of specific exotic mesons but also enhances the systematic classification of hybrid states within the meson nonet. Furthermore, conventional mesons are known to follow Regge trajectories Anisovich:2000kxa . Investigating whether hybrid states also adhere to this pattern is crucial for establishing a complete SU(3) multiplet, as discussed by other authors Chen:2023ukh ; Yan:2023vbh ; Qiu:2022ktc . By addressing these unresolved issues, our work contributes to a deeper understanding of the internal structure of exotic hadronic states and the dynamics of the strong interaction.
@@ III Decay model of the hybrid state
The decay properties of hybrid states are essential for understanding their structure and classification. Several theoretical models have been developed to investigate the decay widths of hybrid states, including the flux tube model Page:1998gz ; Isgur:1985vy ; Close:1994hc ; Barnes:1995hc , the constituent gluon model LeYaouanc:1984gh ; Iddir:1988jd ; Ishida:1991mx ; Kalashnikova:1993xb ; Swanson:1997wy ; Iddir:2000yb ; Ding:2006ya ; Iddir:2007dq ; Benhamida:2019nfx ; Farina:2020slb ; Tanimoto:1982eh ; Tanimoto:1982wy , QCD sum rules DeViron:1984svx ; Zhu:1998sv ; Zhang:2002id ; Chen:2010ic ; Huang:2010dc ; Huang:2016upt , and lattice QCD McNeile:2002az ; McNeile:2006bz ; Woss:2020ayi . In this study, we employ the constituent gluon model, which describes hybrid mesons as quark-antiquark pairs coupled to a transverse electric (TE) gluon with quantum numbers $J_{g}^{PC}=1^{+-}$ Chen:2023ukh . This TE gluon functions as a dynamic degree of freedom, capable of annihilating into a new quark-antiquark pair. The resulting pair then combines with the original quarks to form two mesons, as depicted in Fig. 2. This mechanism provides a comprehensive description of the decay process of hybrid states into meson pairs.
@@
Selection rules are pivotal in determining the decay processes of hybrid mesons. For instance, the spin selection rule prohibits decay into final states composed entirely of spin-singlet mesons if the hybrid’s $q\bar{q}$ configuration is in a spin-singlet state. Additionally, spatial overlap integrals, which describe the overlap of wavefunctions between initial and final states, typically lead to suppressed decay widths for excited states relative to ground states.
Another decay selection rule indicates that, under the assuming $\beta_{B}=\beta_{C}$, the decay of the GE state into $S+S$ is forbidden Iddir:2000yb . This rule arises from the treatment of integrals under the condition $\beta_{B}=\beta_{C}$. In Ref. Iddir:1988jd , the integrals are evaluated with $\beta_{B}\neq\beta_{C}$, and the results indicate that the spatial overlap is generally proportional to $(\beta_{B}-\beta_{C})^{2}$, which is typically small. The same outcome is observed in the flux tube model Close:1994hc . Similarly, if the two S-wave states possess different internal structures or sizes, as discussed in Ref. Close:1994pr , the selection rule prohibiting $S+S$ as final states no longer applies.
@@ IV.1 The $I^{G}J^{PC}=1^{-}1^{-+}$ state
The combined analysis of masses and decay widths provides essential insights into the nature of exotic mesons, particularly those with $J^{PC}=1^{-+}$. Given the quantum numbers $I^{G}J^{PC}=1^{-}1^{-+}$, we assume that $\pi_{1}(1600)$ corresponds to the configuration $(u\bar{u}-d\bar{d})g/\sqrt{2}$ in our calculations. For the decay channels involving $f_{1}(1285)$ and $f_{1}(1420)$, we model them as follows:
@@
Assuming $\pi_{1}(1600)$ as a hybrid state, its total decay width is calculated to be 65 MeV, with $b_{1}(1235)\pi$ identified as the primary decay channel (see Table 3). However, this prediction exhibits a significant discrepancy with Experimental measurements ParticleDataGroup:2024cfk . In the current framework, the $D$-wave contribution to the $b_{1}(1235)\pi$ decay channel is found to vanish, restricting the decay mechanism exclusively to $S$-wave processes. This result is directly conflicts with the PDG-reported branching ratio $\mathcal{B}((b_{1}\pi)_{D\text{-wave}})/\mathcal{B}((b_{1}\pi)_{S\text{-wave}})=0.3\pm 0.1$ ParticleDataGroup:2024cfk . The observed inconsistency challenges the hybrid state interpretation and strongly suggests the need to explore alternative configurations, such as tetraquark states and/or molecular formations.
@@
56.6 8.4 65 $370^{+50}_{-60}$
[…]
@@
where $\alpha_{i}=78.7^{\circ}$ Wang:2019qyy .
From our theoretical framework, the total decay width of $\pi_{1}(2015)$ is determined as 3.21 MeV (Table 4), demonstrating a sharp conflict with the PDG reported upper limits $\Gamma>100$ MeV ParticleDataGroup:2024cfk . The primary decay channel of $\pi_{1}(2015)$ is $b_{1}(1235)\pi$, while other decay channels can be neglected. Experimentally, $\pi_{1}(2015)$ has been observed in the $b_{1}(1235)\pi$ channel. However, the narrow decay width leads us to question the interpretation of $\pi_{1}(2015)$ as the first excited state of the $1^{-}1^{-+}$ hybrid. To date, $\pi_{1}(2015)$ production has been exclusively detected in $b_{1}(1235)\pi$ and $f_{1}(1285)\pi$ final states E852:2004rfa ; E852:2004gpn , with no confirmed observations of hadronic transitions to other channels.
@@
$b_{1}(1235)\pi$ $f_{1}(1285)\pi$ $h_{1}(1170)\rho$ $b_{1}(1235)\omega$ $K_{1}(1270)\bar{K}$
1.75 0.23 $\approx 0$ $\approx 0$ $\approx 0$
$K_{1}(1400)\bar{K}$ $a_{1}(1260)\rho$ $\eta(1295)\pi$ $\eta(1475)\pi$ $\eta_{2}(1645)\pi$
$\approx 0$ $\approx 0$ $0.4$ 0.12 0.3
$\rho(1450)\pi$ $\rho(1700)\pi$ $\pi(1300)\eta$
$310$absent0$\approx$ 0≈ 0 $0.1$
@@
Furthermore, we predict the properties of the first excited isoscalar hybrids $\pi_{1}(2193)$ with $I^{G}J^{PC}=1^{-}1^{-+}$. Its decay properties are listed in Table 5. This state is anticipated to exhibit narrow decay width , with dominant decay channels such as $K_{1}(1650)\bar{K}$ and $b_{1}(1235)\pi$. For the decay channel involving $K_{1}(1650)\bar{K}$, we model $K_{1}(1650)$ as a linear combination of the $2^{1}P_{1}$ and $2^{3}P_{1}$ states, based on meson mass predictions in Ref. Godfrey:1985xj
@@
$b_{1}(1235)\pi$ $f_{1}(1285)\pi$ $h_{1}(1170)\rho$ $b_{1}(1235)\omega$ $K_{1}(1270)\bar{K}$
5.4 0.81 $10$absent0$\approx$ 0≈ 0 0.13
$K_{1}(1400)\bar{K}$ $a_{1}(1260)\rho$ $\eta(1295)\pi$ $\eta(1475)\pi$ $\eta_{2}(1645)\pi$
$\approx 0$ $\approx 0$ $0.86$ 0.37 1.51
$\rho(1450)\pi$ $\rho(1700)\pi$ $\pi(1300)\eta$ $K(1460)\bar{K}$ $\pi(1300)\rho$
$0.9$ 0.25 $0.4$ 0.16 $0.16$
$a_{1}(1260)\eta$ $f_{1}(1420)\pi$ $K_{1}(1270)\bar{K}^{*}$ $K_{1}(1650)\bar{K}$
0.2 0.12 $\approx 0$ $20.1$